OpenStudy (anonymous):

jhon has a debt of $35000 the payment day is in 16 months,he pretend to pay $12000 at the end of 158 days,and $13000 189 days after the first payment,how much he has to pay in payment day to discharge from debt?the interest rate is 50% in a year.(simple interest)

OpenStudy (anonymous):

the answer in book is $2231.95

OpenStudy (experimentx):

50% in year .. 50/360% per day ... perhaps this is step problem.

OpenStudy (anonymous):

my ans is 28715.69

OpenStudy (experimentx):

a = 35k + 35k*0.5/360 * 158 - 12k b = a + a*0.5/360 * 189 - 13k c = b + a*0.5/360 * (16*30 - 159 - 158) ans

OpenStudy (anonymous):

32679.9

OpenStudy (anonymous):

i think the answ is wrong too.

OpenStudy (anonymous):

for me that $2231.95 is impossible

OpenStudy (anonymous):

but in second line you did not use simple interest

OpenStudy (experimentx):

oh ... on which line?

OpenStudy (anonymous):

when you calculate b

OpenStudy (experimentx):

oh ... that's because you already paid off 12k ... if you have to pay again interest of 12k then it would be unfair .... is it like fixed deposit loan or what??

OpenStudy (anonymous):